OpenStudy (anonymous):

Tangent line equation. (a, f(a)), f(x) = x^2, with a = 2.

OpenStudy (anonymous):

there are 2 ways to solve this problem

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

1) By using the definition and I forgot abt the other way lol

OpenStudy (anonymous):

hahaha , ok, lets use the definition

OpenStudy (anonymous):

first : (a, f(a) ) is going to be : (2, 4)<-- you'll use these points later to find the equation we'll have to find the slope first

OpenStudy (anonymous):

>_< calculus I lol, alright, remind me, m = limit of that function

OpenStudy (anonymous):

f(x) - f(a) = f'(a) . (x - a)

OpenStudy (anonymous):

LOL! that's it! find the derivative!

OpenStudy (anonymous):

you have f(a) = 4, and a = 2

OpenStudy (anonymous):

are you sure that you have to find the derivative? I remember something with limits

OpenStudy (anonymous):

\[\ m= lim_{h \rightarrow 0}f(a) - f(a+1)/h\] that's the equation needed if we want to find the slope using the definition , right?

OpenStudy (anonymous):

yeap!

OpenStudy (anonymous):

alright! all we have to do is plug in the values we have

OpenStudy (anonymous):

hmmm if f(x) = x^2, then f(a) = a^2 , right?

OpenStudy (anonymous):

that's correct.

OpenStudy (anonymous):

then,\[= \lim_{h \rightarrow 0} (a^2 -a^2-1)/h\] = -1/h...hmm there's something wrong. h must be canceled

OpenStudy (anonymous):

LOL! no no no wait!

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} f(a) - f(a+h)/h \] that's the right equation LOL! sorry ^^"

OpenStudy (anonymous):

np

OpenStudy (anonymous):

so you'll get:\[=\lim_{h \rightarrow 0} (a^2 -a^2 -h)/h\]\[=\lim_{h \rightarrow 0} -h/h\] m = -1 ^_^ there we have found m and we have the points (2,4) <-- you know how I got them right?

OpenStudy (anonymous):

a = 2, f(a) = (2)^2 = 4 (2,4)

OpenStudy (anonymous):

ahhhh

OpenStudy (anonymous):

i got it. thanks!

OpenStudy (anonymous):

soooooooooooo, we'll use the equation which is : y -yo = m (x-xo) y - 4 = = -1(x-2) ^_^ LOL np :)

OpenStudy (anonymous):

sorry abt the mess though >_<

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

^_^ good luck